Please solve following Concrete Reinforcement question ( ACI318-14 ) to be used.

3) fe - 3000 Psi fue 6o kei FIND 0 Mn CHECK E+ O R 2-#7 24 3-#11 14
Answer #1

V D med = 2.5 287 d = 26.5 3 #11 2.5 b=14 wildth (b) = 4 =1911 Elective depth (d ) = (24+2,5) =26.5 11 = 2.5 d As = 2To find out strains in compression steel to verity the assumption that it is yielding: we know that: 6,88 -2.5 3(ed)= ( )x0.→ Calculation of strains, stresses and steel arecco 7.03-25 know that & = cod! -205) Ox0-00 c C-d x0·003 - 7.03 C 0.00 1933

answered by: ANURANJAN SARSAM
Answer #2

ANSWER :V D med = 2.5 287 d = 26.5 3 #11 2.5 b=14 wildth (b) = 4 =1911 Elective depth (d ) = (24+2,5) =26.5 11 = 2.5 d As = 2To find out strains in compression steel to verity the assumption that it is yielding: we know that: 6,88 -2.5 3(ed)= ( )x0.→ Calculation of strains, stresses and steel arecco 7.03-25 know that & = cod! -205) Ox0-00 c C-d x0·003 - 7.03 C 0.00 1933